# Magnetic Anti-de Sitter Wormholes as seeds for Higgs Inflation

### v1

@@ I Introduction
The tremendous advances of theoretical and experimental physics during the previous century led to an extremely detailed understanding of the history and evolution of our Universe, from the present era back to its primordial origin, when its size was billion times smaller than that of a proton. Cosmic inflation Kazanas:1980tx ; Guth:1980zm ; Linde:1981mu is our leading candidate theory describing the physics of the rapidly expanding early universe at this primordial epoch, leading to predictions that conform very well with our current cosmological observational data - a surprising feat considering the vast hierarchy of scales between the primordial and current epoch.
@@
Nevertheless the onset of inflation and the epochs preceding it were never really understood in detail, despite numerous bold efforts of theoretical physicists such as Hartle:1983ai ; Vilenkin:1983xq . Effective field theory itself offers an operationally concrete setting to pose the relevant questions: Why and how did the universe start with appropriate conditions high up on the inflaton potential? While these being initial/boundary value questions within the context of purely inflationary theory, we do expect that a more complete theory could explain them in a meaningful and natural way, and perhaps in a more dynamical fashion.
@@
In this Letter we propose and show how certain types of semi-classical solutions of General Relativity coupled to the Standard Model can catalyze the onset of inflation. Our paradigm is based on two assumptions: that the inflaton is to be identified with the experimentally observed Higgs particle (the so called “Higgs inflation” scenario Bezrukov:2007ep ), and that the effective Higgs potential turns negative at a certain high energy window Bezrukov:2012sa ; Buttazzo:2013uya . While the first assumption is not entirely crucial for the viability of our scenario, it naturally resolves the problem of explaining what happens to the inflaton in our current energy scales and why we have not yet observed any other scalar particle besides the Higgs. The second assumption while being structurally more important for our proposal, it aligns with current experimental data for the Higgs and Top quark masses ParticleDataGroup:2024cfk that lead to a metastability of the electroweak vacuum Buttazzo:2013uya .
@@
Our proposed model has various phenomenologically interesting outputs. As other models of Higgs inflation (that essentially generalise Starobinsky’s $R^{2}$ model Starobinsky:1980te ), it can explain very well current cosmological data Planck:2018jri ; BICEP:2021xfz . At the same time it fares better regarding some troublesome points of former Higgs inflation models such as the unitarity issue Barbon:2009ya ; Bezrukov:2010jz . It also predicts the presence of primordial magnetic fields with a very large coherence length. While these magnetic fields get diluted to an enormous degree due to the accelerated inflationary expansion Turner:1987bw , they could perhaps play an important role even at later stages of the evolution of the universe if further physical mechanisms that can amplify and sustain them become active Hogan:1983zz ; Turner:1987bw ; Kandus:2010nw ; Durrer:2013pga ; Subramanian:2015lua . Finally, it has the potential of being embeddable in a UV complete model of quantum gravity such as string theory, due to the Euclidean AdS asymptotics of the relevant wormhole backgrounds Maldacena:2004rf ; Betzios:2019rds ; Betzios:2021fnm ; VanRaamsdonk:2021qgv ; Antonini:2022ptt ; Betzios:2024oli .
@@ II The Model
where $\kappa\equiv 1/M_{\rm Pl}^{2}$ and $V(\phi)$ is the scalar field potential. The non-minimal coupling term typically arises, when considering the one-loop effective action of the SM coupled to gravity via quantum loop corrections and is required to renormalise the stress energy tensor of scalar fields in curved backgrounds Birrell:1982ix . The scalar potential, should then also be understood to be the loop-corrected effective potential. In general, the EFT contains higher derivative and more complicated non-minimal coupling terms, such as $W(\phi)R^{m}$. In this work we focus in the minimal Higgs inflation scenario where the $\xi\phi^{2}R$ term is the most relevant term in such an expansion Bezrukov:2007ep ; Bezrukov:2010jz . This is further corroborated by the fact that the characteristic size of the throat of our wormhole backgrounds, is at least $10^{5}M_{\rm Pl}^{-1}$, rendering higher curvature and quantum corrections subdominant [^1] See the Supplementary Material for details..
@@
$\mathcal{S}_{E}=\int{\rm d}^{4}x\sqrt{g_{E}}\left(-\frac{R}{2\kappa}+\frac{\partial_{\mu}\chi\partial^{\mu}\chi}{2}+U(\chi)+\mathcal{L}^{E}_{\rm rad,matter}\right)$ (4)
@@
Since we want to understand the dynamics of the early universe (having a small size), the leading contribution to the Einstein equations stems from the stress energy tensor of possible axionic fields scaling as $a^{-6}$, followed by that of the radiation and (non-relativistic) matter fields, scaling as $a^{-4}$ and $a^{-3}$ respectively ($a$ is the scale factor in the FLRW ansatze (7)). The axionic contribution was analysed in Betzios:2024oli , but in what follows, we shall only consider the radiation (density and pressure) components to the stress energy tensor $\rho_{\rm rad},P_{\rm rad}$, arising from the experimentally observed SM gauge fields [^2] We also include in $\rho_{\rm rad}$ the contribution of any relativistic matter such as neutrinos. To first approximation we neglect any non-relativistic matter contribution since it is subdominant.. Another advantage is that their actions are scale-invariant and so their contribution is also invariant under the rescaling (2).
Our approach is phenomenological, as we do not solve the resulting equations of motion (EOMs) for the radiation fields directly, but treat them as a fluid (see Rey:1989th ; Hosoya:1989zn ; Betzios:2019rds ; Marolf:2021kjc ; Lan:2024gnv for wormhole examples with radiation fields where the complete field equations are solved). However, this approach suffices to demonstrate the existence of the backgrounds we are interested in and has the merit that it applies to radiation fields in a universal way – all the SM radiation fields contribute to our background solutions. While in their Lorentzian equation of state
@@
their energy density is positive definite, the “Euclidean energy” density of the radiation, $\rho_{\rm rad}^{E}$, can be negative if the field has a dominant magnetic component. For example, in the $U(1)$ E/M case the “energy” density from the Euclidean stress energy tensor is Gupta:1989bs
@@
The observed isotropy and homogeneity of our Universe, leads us to use the following Euclidean FLRW type of ansatze
@@
where the prime denotes a derivative with respect to $\tau$. The conditions that describe a “wineglass” shaped wormhole are Betzios:2024oli
$a^{\prime\prime}(0)<0\,,\quad a^{\prime}(0)=0\,,\quad\chi^{\prime}(0)=0\,,$ (9)
where the scale factor exhibits a local maximum, $a(0)=a_{\rm max}$. These were shown to be good initial conditions for a subsequent Lorentzian inflationary evolution after the Universe has nucleated Betzios:2024oli . Additionally, there is another point in the Euclidean evolution where the scale factor must reach a local minimum, $a_{\rm min}$, i.e. $a^{\prime}(\tau_{\rm min})=0$ and $a^{\prime\prime}(\tau_{\rm min})>0$. Since the Euclidean evolution of the scalar happens in the potential $-\tilde{U}(\chi)$ Betzios:2024oli , in order to find the wormhole solutions of our interest, we assume an effective potential $\tilde{U}(\chi)$ with a positive global maximum and a negative global minimum, for which (see fig. 1 for an example)
@@
A short analysis shows that it is impossible to satisfy the wormhole conditions when $\tilde{\rho}_{\rm rad}^{E}>0$.
@@
so that the scale factor is real. We now find both branches of solutions $x_{\pm}$, $x_{+}\geq x_{-}$, becoming the same at the boundary of the inequality above, see fig. 2. We therefore choose $a_{\tau_{\rm min}}$ in the $x_{-}$ and $a_{max}$ in the $x_{+}$ branch. For the $a_{max}$ solution ($\mathcal{H}_{E}=-\tilde{U}_{0}$) we find that $a^{\prime\prime}(0)<0$ always. For the $x_{-}$ solution, we find that $a^{\prime\prime}(\tau_{\rm min})>0$ as long as $\mathcal{H}_{E}^{\rm max}>\mathcal{H}_{E}(\tau_{\rm min})>-\tilde{U}_{0}$. The existence of the upper bound is tied to the fact that if $\mathcal{H}_{E}$ is too high at $\tau_{\rm min}$ (for example if one sends $\tilde{U}_{\text{global min.}}\rightarrow-\infty$ in fig. 1), the friction will not be enough to stop the particle, resulting in overshooting. In the rest we shall assume that the effective potential does not exhibit such pathological characteristics [^4] Another possibility is that the potential in the Euclidean region has both a local minimum and maximum as studied in Betzios:2024oli , so that the Euclidean motion in the inverted potential cannot exhibit overshooting..
The result of this analysis is that we can generically obtain the wanted wormhole solutions, using a predominently magnetic type of radiation for a spatially compact early universe, as long as the inequality (13) is respected. In this Letter we do not analyse further possible mechanisms by which such primordial magnetic fields can be produced, see Giovannini:2003yn ; Brandenburg:2004jv ; Barrow:2006ch ; Subramanian:2015lua ; Kandus:2010nw for reviews that focus on their generation during and after inflation.
@@ III Higgs inflation
As discussed earlier, our model can be embedded in Higgs inflation Bezrukov:2007ep , if we consider a tree level potential $V(\phi)=\lambda(\phi^{2}-v_{\rm ew}^{2})^{2}/4$, where quantum loop effects induce a running quartic coupling, $\lambda(\mu)$ ($\mu$ is the RG scale), that remains positive up to the inflationary scale $\mu_{\star}$. The non-minimal coupling, $\xi(\mu)$, has the effect of flattening the effective potential at large field values in the Einstein frame, which can lead to a sufficient number of $e-$ folds for inflation. Assuming values for the Higgs and Top quark mass around the central values provided by ParticleDataGroup:2024cfk , we seem to be in the metastable region for the EW vacuum. This introduces a negative minimum ($v_{\rm true}$) in the effective potential, and the quartic coupling turns negative in a certain energy window. We are interested in the parametric regime which places this window above the inflationary scale $\mu_{\star}$, corresponding to earlier times in the evolution of the universe prior to the onset of inflation [^5] In Bezrukov:2014ipa , an alternative scenario was considered in which the true (negative) minimum lies between the electroweak vacuum and the plateau region and hence affects the Lorentzian time evolution. This scenario is not very satisfactory, since unless one invokes certain thermal effects that completely lift this minimum, inflation will end there and the resulting cosmology will be in stark contrast with observations.. In this scenario, we obtain an inflationary hilltop plateau for $v_{\rm true}>\chi>v_{\rm ew}$, and the shape of the quantum corrected effective potential is shown schematically in the upper panel of Figure 1. While we obtain this scenario for appropriate values of SM parameters, well within experimental bounds (see Fumagalli:2016lls ; Enckell:2018kkc ; Masina:2024ybn for more detailed discussions and a list of references), we remain agnostic regarding the actual value of $v_{\rm true}$ and the detailed shape of the effective potential at such high energy scales, since it is sensitive to the precise UV-completion of our model and the scale at which beyond the SM physics, such as supersymmetry, could kick in. Nevertheless in low energy effective actions arising from string theory, one generically finds effective potentials with a global negative minimum, typically preserving some amount of supersymmetry, so our scenario is very natural in potential string motivated UV completions of our model [^6] In this case, one finds that the opposite is harder to achieve - that is to uplift a part of the effective scalar potential to positive values in a controllable way, see Baumann:2014nda for a review..
@@
If we adopt our scenario of Higgs inflation, the potential in the plateau region at horizon crossing is approximately $U(\chi_{\star})\simeq\lambda_{\star}/(4\kappa^{2}\xi^{2}_{\star})$ and this ratio of the couplings remains approximately constant for most of the inflationary period due to the slow-roll condition. The amplitude of the primordial curvature perturbation on super-Hubble scales generated by single-field slow-roll inflation, $A^{\star}_{s}=\frac{N_{\star}^{2}\lambda_{\star}}{72\pi^{2}\xi^{2}_{\star}}$, is constrained by recent observations Planck:2018jri to approximately $\sim 2.1\times 10^{-9}$ at the pivot scale $k_{\star}=0.05$ Mpc-1. This constraint, in turn, determines the ratio $\lambda_{\star}/\xi^{2}_{\star}\sim(4-5)\times 10^{-10}$ and $U(\chi_{\star})\simeq(1.0-1.2)\times 10^{-10}\,M_{\rm Pl}^{4}$, for $N_{\star}=60-55$ $e$-folds of inflation.
@@
We should emphasize at this point that these conditions are at the inflationary scale, at lower energies from the end of inflation to the current EW scale $\lambda(\mu)$ has an appreciable non-trivial running, so that Buttazzo:2013uya $\lambda_{\rm ew}\simeq 0.126\gg\lambda_{\star}$ (above the inflationary scale it turns negative). Hence our model accommodates a non-minimal coupling whose value $\xi_{\star}$ at the inflationary scale is much lower than that of the standard tree level discussions of Higgs inflation, where $\lambda_{\star}\simeq\lambda_{\rm ew}$ Bezrukov:2007ep ; Rubio:2018ogq . This result helps with the issue of violating perturbative unitarity at the scale $\Lambda_{u}=M_{\rm Pl}/\xi$, that Higgs inflation models exhibit Burgess:2009ea ; Barbon:2009ya ; Burgess:2010zq ; Bezrukov:2010jz , since for small $\xi_{\star}$ it is pushed well above the inflationary scale, and towards the Planck scale, where we inevitably need to assume a certain UV-completion of our EFT model.
@@
and
${\rho_{\rm rad}^{E}}_{\star}\simeq\frac{\kappa{U}_{0}^{2}\tilde{\rho}^{E}_{\rm rad}}{3}\,,\quad 0>{\rho_{\rm rad}^{E}}_{\star}>-(2.5-3.0)\times 10^{-11}M_{\rm Pl}^{4}\,,$ (15)
were we used eqns. (13) and (14). We then separate from this density its magnetic component ${\rho_{\rm rad}^{E}}_{\star}={\rho_{\rm rest}^{E}}_{\star}-B^{2}_{\star}/2$. Assuming that no fine-tuning takes place, the scale of the magnetic field density at the beginning of inflation is $B^{2}_{\star}\simeq 10^{-11}M_{\rm Pl}^{4}$. During inflation, the universe was a very poor conductor, leading to an adiabatic dilution of this primordial density by a factor of $\simeq 10^{-96}-10^{-104}$ for $N_{\star}=55-60$ $e$-folds, resulting into $B_{\rm reh}\simeq 1000-0.1\,{\rm G}$ at the beginning of reheating. The rapid expansion also leads to fields with very large coherence length after horizon re-entry. During reheating and up until horizon re-entry the universe turns into a very good conductor. By the time we enter the radiation era, the electric fields get eliminated by the currents of the highly conducting plasma, and the universe is left with classical large-scale magnetic fields Turner:1987bw . One can then define a ratio $r_{B}={B^{2}}/8\pi{\rho_{\gamma}}$ ($\rho_{\gamma}$ is the energy density of relativistic particle species), that remains approximately constant during the history of the universe from reheating until galaxy formation Turner:1987bw . Possible galactic dynamo mechanisms Kulsrud:1992rk ; Beck:1995zs ; Kulsrud:1999bg ; Widrow:2002ud ; Brandenburg:2004jv that can sustain the currently observed magnetic fields require a minimum “seed” field of the order of $B_{\rm seed}\simeq 10^{-22}$ G (that is model dependent) and a coherence scale of at least $10$ Kpc. Since $\rho_{\gamma}\simeq 10^{-51}\,{\rm GeV}^{4}$ today, we find a sufficient seed ratio $r_{B}\simeq 10^{-34}$. Evolving back the current radiation fields to the reheating era, we find $\rho_{\gamma,{\rm reh}}\simeq 10^{62}\,{\rm GeV}^{4}$ assuming instantaneous thermalization. In consistency with Turner:1987bw we find that the magnetic fields in our model after reheating, while being coherent to extremely large scales, they get diluted to such a high degree during the expansion of the universe, making it impossible for them to seed the galactic dynamo mechanism unless additional physical mechanisms are invoked Hogan:1983zz ; Turner:1987bw ; Kandus:2010nw ; Subramanian:2015lua .

### v2

@@ I Introduction
The tremendous advances of theoretical and experimental physics during the previous century led to an extremely detailed understanding of the history and evolution of our Universe, from the present era back to its primordial origin, when its size was billion times smaller than that of a proton. Cosmic inflation Kazanas (1980); Guth (1981); Linde (1982) is our leading candidate theory describing the physics of the rapidly expanding early universe at this primordial epoch, leading to predictions that conform very well with our current cosmological observational data - a surprising feat considering the vast hierarchy of scales between the primordial and current epoch.
@@
Nevertheless the onset of inflation and the epochs preceding it were never really understood in detail, despite numerous bold efforts of theoretical physicists such as Hartle and Hawking (1983); Vilenkin (1983). Effective field theory itself offers an operationally concrete setting to pose the relevant questions: Why and how did the universe start with appropriate conditions high up on the inflaton potential? While these being initial/boundary value questions within the context of purely inflationary theory, we do expect that a more complete theory could explain them in a meaningful and natural way, and perhaps in a more dynamical fashion.
In this paper we propose and show how certain types of semi-classical solutions of General Relativity coupled to the Standard Model can catalyze the onset of inflation. These are $Z_{2}$ ($\tau\leftrightarrow-\tau$) symmetric “wineglass anti-de Sitter wormholes” Betzios and Papadoulaki (2024) that describe the non-perturbative nucleation process of the inflationary universe, whose initial state is determined from the $\tau=0$ slice of the Euclidean wormhole geometry.
@@
Our paradigm is based on two assumptions: that the inflaton is to be identified with the experimentally observed Higgs particle (the so called “Higgs inflation” scenario Bezrukov and Shaposhnikov (2008)), and that the effective Higgs potential turns negative at a certain high energy window Bezrukov et al. (2012); Buttazzo et al. (2013). While the first assumption is not entirely crucial for the viability of our scenario, it naturally resolves the problem of explaining what happens to the inflaton in our current energy scales and why we have not yet observed any other scalar particle besides the Higgs. The second assumption while being structurally more important for our proposal, it aligns with current experimental data for the Higgs and Top quark masses Navas et al. (2024) that lead to a metastability of the electroweak vacuum Buttazzo et al. (2013).
@@
Our proposed model has various phenomenologically interesting outputs. As other models of Higgs inflation (that essentially generalise Starobinsky’s $R^{2}$ model Starobinsky (1980)), it can explain very well current cosmological data Akrami et al. (2020); Ade et al. (2021). At the same time it fares better regarding some troublesome points of former Higgs inflation models such as the unitarity issue Barbon and Espinosa (2009); Bezrukov et al. (2011). It also predicts the presence of primordial magnetic fields with a very large coherence length. While these magnetic fields get diluted to an enormous degree due to the accelerated inflationary expansion Turner and Widrow (1988), they could perhaps play an important role even at later stages of the evolution of the universe if further physical mechanisms that can amplify and sustain them become active Hogan (1983); Turner and Widrow (1988); Kandus et al. (2011); Durrer and Neronov (2013); Subramanian (2016). Finally, it has the potential of being embeddable in a UV complete model of quantum gravity such as string theory, due to the Euclidean AdS asymptotics of the relevant wormhole backgrounds Maldacena and Maoz (2004); Betzios et al. (2019, 2022); Van Raamsdonk (2022); Antonini et al. (2023); Betzios and Papadoulaki (2024).
@@ II The Model
where $\kappa\equiv 1/M_{\rm Pl}^{2}$ and $V(\phi)$ is the scalar field potential. The non-minimal coupling term typically arises, when considering the one-loop effective action of the SM coupled to gravity via quantum loop corrections and is required to renormalise the stress energy tensor of scalar fields in curved backgrounds Birrell and Davies (1982). The scalar potential, should then also be understood to be the loop-corrected effective potential. In general, the EFT contains higher derivative and more complicated non-minimal coupling terms, such as $W(\phi)R^{m}$. In this work we focus in the minimal Higgs inflation scenario where the $\xi\phi^{2}R$ term is the most relevant term in such an expansion Bezrukov and Shaposhnikov (2008); Bezrukov et al. (2011). This is further corroborated by the fact that the characteristic size of the throat of our wormhole backgrounds, is at least $10^{5}M_{\rm Pl}^{-1}$, rendering higher curvature and quantum corrections subdominant  [^1] See the Supplementary Material for details..
@@
$\mathcal{S}_{E}=\int{\rm d}^{4}x\sqrt{g_{E}}\left(-\frac{R}{2\kappa}+\frac{\partial_{\mu}\chi\partial^{\mu}\chi}{2}+U(\chi)+\mathcal{L}^{E}_{\rm rad,matter}\right)\,,$ (4)
@@
Since we want to understand the dynamics of the early universe (having a small size), the leading contribution to the Einstein equations stems from the stress energy tensor of possible axionic fields scaling as $a^{-6}$, followed by that of the radiation and (non-relativistic) matter fields, scaling as $a^{-4}$ and $a^{-3}$ respectively ($a$ is the scale factor in the FLRW ansatze (7)). The axionic contribution was analysed in Betzios and Papadoulaki (2024), but in what follows, we shall only consider the radiation (density and pressure) components to the stress energy tensor $\rho_{\rm rad},P_{\rm rad}$, arising from the experimentally observed SM gauge fields  [^2] We also include in $\rho_{\rm rad}$ the contribution of any relativistic matter such as neutrinos. To first approximation we neglect any non-relativistic matter contribution since it is subdominant.. Another advantage is that their actions are scale-invariant and so their contribution is also invariant under the rescaling (2).
Our approach is phenomenological, as we do not solve the resulting equations of motion (EOMs) for the radiation fields directly, but treat them as a fluid (see Rey (1990); Hosoya and Ogura (1989); Betzios et al. (2019); Marolf and Santos (2021); Deshpande and Lunin (2023); Lan and Piao (2024) for wormhole examples with radiation fields where the complete field equations are solved). However, this approach suffices to demonstrate the existence of the backgrounds we are interested in and has the merit that it applies to radiation fields in a universal way – all the SM radiation fields contribute to our background solutions. While in their Lorentzian equation of state
@@
their energy density is positive definite, the “Euclidean energy” density of the radiation, $\rho_{\rm rad}^{E}$, can be negative if the field has a dominant magnetic component. For example, in the $U(1)$ E/M case the “energy” density from the Euclidean stress energy tensor is Gupta et al. (1990)
@@
Arguments regarding the dominance of $SO(4)$ symmetric Euclidean configurations in the path integral  [^3] See for example the lectures Coleman (1985)., consistent with the observed isotropy and homogeneity of our universe, lead us to use the following Euclidean FLRW type of ansatze for describing its nucleation process
@@
where the prime denotes a derivative with respect to $\tau$. The conditions that describe a “wineglass” shaped wormhole with a local maximum $a(0)=a_{\rm max}$ for the scale factor are Betzios and Papadoulaki (2024)
$a^{\prime\prime}(0)<0\,,\quad a^{\prime}(0)=0\,,\quad\chi^{\prime}(0)=0\,.$ (9)
These are favorable initial conditions for a subsequent Lorentzian inflationary evolution after the universe nucleates at $\tau=0$ Betzios and Papadoulaki (2024), since upon $\tau=it,\,$ ${\rm d}^{2}a/{\rm d}t^{2}(0)>0$. Additionally, there is another point in the Euclidean evolution where the scale factor must reach a local minimum, $a_{\rm min}$, i.e. $a^{\prime}(\tau_{\rm min})=0$ and $a^{\prime\prime}(\tau_{\rm min})>0$. Since the Euclidean evolution of the scalar happens in the potential $-\tilde{U}(\chi)$ Betzios and Papadoulaki (2024), in order to find the wormhole solutions of our interest, we assume an effective potential $\tilde{U}(\chi)$ with a positive global maximum and a negative global minimum, for which (see fig. 1 for an example)
@@
A short analysis shows that it is impossible to satisfy the wormhole conditions when $\tilde{\rho}_{\rm rad}^{E}>0$, so a negative energy density is needed. This requirement arises from a Euclidean analog of gravitational focusing where gravity shrinks the space/scale factor, while negative Euclidean energy (magnetic flux in our case) enables a bounce and re-expansion. The competition is between the curvature term $1/a^{2}$ and the flux term $\sim 1/a^{4}$, which dominates at small $a$. Other fields, like axions ($\sim 1/a^{6}$) or magnetized matter ($\sim 1/a^{3}$), can play a similar role.
@@
so that the scale factor is real. We now find both branches of solutions $x_{\pm}$, $x_{+}\geq x_{-}$, becoming the same at the boundary of the inequality above, see fig. 2. We therefore choose $a_{\tau_{\rm min}}$ in the $x_{-}$ and $a_{max}$ in the $x_{+}$ branch. For the $a_{max}$ solution ($\mathcal{H}_{E}=-\tilde{U}_{0}$) we find that $a^{\prime\prime}(0)<0$ always. For the $x_{-}$ solution, we find that $a^{\prime\prime}(\tau_{\rm min})>0$ as long as $\mathcal{H}_{E}^{\rm max}>\mathcal{H}_{E}(\tau_{\rm min})>-\tilde{U}_{0}$. The existence of the upper bound is tied to the fact that if $\mathcal{H}_{E}$ is too high at $\tau_{\rm min}$ (for example if one sends $\tilde{U}_{\text{global min.}}\rightarrow-\infty$ in fig. 1), the friction will not be enough to stop the particle, resulting in overshooting. In the rest we shall assume that the effective potential does not exhibit such pathological characteristics  [^5] Another possibility is that the potential in the Euclidean region has both a local minimum and maximum as studied in Betzios and Papadoulaki (2024), so that the Euclidean motion in the inverted potential cannot exhibit overshooting..
The result of this analysis is that we can generically obtain the wanted wormhole solutions for a spatially compact early universe, using relativistic fields with a slightly dominant magnetic component, as long as the inequality (13) is respected. In this paper we do not analyse further possible mechanisms by which such primordial magnetic fields can be produced, see Giovannini (2004); Brandenburg and Subramanian (2005); Barrow et al. (2007); Subramanian (2016); Kandus et al. (2011) for reviews that focus on their generation during and after inflation.
@@ III Higgs inflation
As discussed earlier, our model can be embedded in Higgs inflation Bezrukov and Shaposhnikov (2008), if we consider a tree level potential $V(\phi)=\lambda(\phi^{2}-v_{\rm ew}^{2})^{2}/4$, where quantum loop effects induce a running quartic coupling, $\lambda(\mu)$ ($\mu$ is the renormalization group (RG) scale), that remains positive up to the inflationary scale $\mu_{\star}$. The non-minimal coupling, $\xi(\mu)$, has the effect of flattening the effective potential at large field values in the Einstein frame, which can lead to a sufficient number of $e-$ folds for inflation. Assuming values for the Higgs and Top quark mass around the central values provided by Navas et al. (2024), we seem to be in the metastable region for the EW vacuum. This introduces a negative minimum ($v_{\rm true}$) in the effective potential, and the quartic coupling turns negative in a certain energy window  [^6] If the Higgs mass, the SU(3) gauge coupling, and other measured SM parameters are fixed at their central values, there is always a critical top quark mass for which the electroweak and Planck-scale vacua become degenerate (see, e.g. Masina and Quiros (2024)). Below this threshold, the Higgs effective potential remains stable, while for values above it, becomes metastable. Experimental data place us in the metastable region Navas et al. (2024), supporting our proposal. While other parameters ($m_{h},g_{3},M_{Z}$, etc.) affect stability, a full analysis of the parameter space is beyond the scope of this paper.. We are interested in the parametric regime which places this window above the inflationary scale $\mu_{\star}$, corresponding to earlier times in the evolution of the universe prior to the onset of inflation  [^7] In Bezrukov et al. (2015), an alternative scenario was considered in which the true (negative) minimum lies between the electroweak vacuum and the plateau region and hence affects the Lorentzian time evolution. This scenario is not very satisfactory, since unless one invokes certain thermal effects that completely lift this minimum, inflation will end there and the resulting cosmology will be in stark contrast with observations.. In this scenario, we obtain an inflationary hilltop plateau for $v_{\rm true}>\chi>v_{\rm ew}$, and the shape of the quantum corrected effective potential is shown schematically in the upper panel of Figure 1. While we obtain this scenario for appropriate values of SM parameters, well within experimental bounds (see Fumagalli and Postma (2016); Enckell et al. (2018); Masina and Quiros (2024) for more detailed discussions and a list of references), we remain agnostic regarding the actual value of $v_{\rm true}$ and the detailed shape of the effective potential at such high energy scales, since it is sensitive to the precise UV-completion of our model and the scale at which beyond the SM physics, such as supersymmetry, could kick in. Nevertheless in low energy effective actions arising from string theory, one generically finds effective potentials with a global negative minimum, typically preserving some amount of supersymmetry, so our scenario is very natural in potential string motivated UV completions of our model  [^8] In this case, one finds that the opposite is harder to achieve - that is to uplift a part of the effective scalar potential to positive values in a controllable way, see Baumann and McAllister (2015) for a review..
@@
If we adopt our scenario of Higgs inflation, the potential in the plateau region at horizon crossing is approximately $U(\chi_{\star})\simeq\lambda_{\star}/(4\kappa^{2}\xi^{2}_{\star})$ and this ratio of the couplings remains approximately constant for most of the inflationary period due to the slow-roll condition. The amplitude of the primordial curvature perturbation on super-Hubble scales generated by single-field slow-roll inflation, $A^{\star}_{s}=\frac{N_{\star}^{2}\lambda_{\star}}{72\pi^{2}\xi^{2}_{\star}}$, is constrained by recent observations Akrami et al. (2020) to approximately $\sim 2.1\times 10^{-9}$ at the pivot scale $k_{\star}=0.05$ Mpc-1. This constraint, in turn, determines the ratio $\lambda_{\star}/\xi^{2}_{\star}\sim(4-5)\times 10^{-10}$ and $U(\chi_{\star})\simeq(1.0-1.2)\times 10^{-10}\,M_{\rm Pl}^{4}$, for $N_{\star}=60-55$ $e$-folds of inflation.
@@
We should emphasize at this point that these conditions are at the inflationary scale, at lower energies from the end of inflation to the current EW scale $\lambda(\mu)$ has an appreciable non-trivial running, so that Buttazzo et al. (2013) $\lambda_{\rm ew}\simeq 0.126\gg\lambda_{\star}$ (above the inflationary scale it turns negative). Hence our model accommodates a non-minimal coupling whose value $\xi_{\star}$ at the inflationary scale is much lower than that of the standard tree level discussions of Higgs inflation, where $\lambda_{\star}\simeq\lambda_{\rm ew}$ Bezrukov and Shaposhnikov (2008); Rubio (2019). This result helps with the issue of violating perturbative unitarity at the scale $\Lambda_{u}=M_{\rm Pl}/\xi$, that Higgs inflation models exhibit Burgess et al. (2009); Barbon and Espinosa (2009); Burgess et al. (2010); Bezrukov et al. (2011), since for small $\xi_{\star}$ it is pushed well above the inflationary scale, and towards the Planck scale, where we inevitably need to assume a certain UV-completion of our EFT model.
@@
Using eqns. (11) and (13) we find that the maximum value of the scale factor is approximately $a^{2}_{\rm max}\simeq{3}/{(\kappa U_{0})}$. Substituting this in the definition $\kappa\rho_{\rm rad}^{E}/3=\tilde{\rho}_{\rm rad}^{E}/a^{4}$ and using eq. (14) we obtain
${\rho_{\rm rad}^{E}}_{\star}\simeq\frac{\kappa{U}_{0}^{2}\tilde{\rho}^{E}_{\rm rad}}{3}\,,\quad 0>{\rho_{\rm rad}^{E}}_{\star}>-(2.5-3.0)\times 10^{-11}M_{\rm Pl}^{4}\,.$ (15)
We then separate from this density its magnetic component ${\rho_{\rm rad}^{E}}_{\star}={\rho_{\rm rest}^{E}}_{\star}-B^{2}_{\star}/2$. Assuming that no fine-tuning takes place, but rather a small asymmetry between the magnitudes of the magnetic and electric components, the scale of the magnetic field density at the beginning of inflation is $B^{2}_{\star}\simeq 10^{-11}M_{\rm Pl}^{4}$. During inflation, the universe was a very poor conductor, leading to an adiabatic dilution of this primordial density by a factor of $\simeq 10^{-96}-10^{-104}$ for $N_{\star}=55-60$ $e$-folds, resulting into $B_{\rm reh}\simeq 1000-0.1\,{\rm G}$ at the beginning of reheating. The rapid expansion also leads to fields with very large coherence length after horizon re-entry. During reheating and up until horizon re-entry the universe turns into a very good conductor. By the time we enter the radiation era, the electric fields get eliminated by the currents of the highly conducting plasma, and the universe is left with classical large-scale magnetic fields Turner and Widrow (1988). One can then define a ratio $r_{B}={B^{2}}/8\pi{\rho_{\gamma}}$ ($\rho_{\gamma}$ is the energy density of relativistic particle species), that remains approximately constant during the history of the universe from reheating until galaxy formation Turner and Widrow (1988). Possible galactic dynamo mechanisms Kulsrud and Anderson (1992); Beck et al. (1996); Kulsrud (1999); Widrow (2002); Brandenburg and Subramanian (2005) that can sustain the currently observed magnetic fields require a minimum “seed” field of the order of $B_{\rm seed}\simeq 10^{-22}$ G (that is model dependent) and a coherence scale of at least $10$ Kpc. Since $\rho_{\gamma}\simeq 10^{-51}\,{\rm GeV}^{4}$ today, we find a sufficient seed ratio $r_{B}\simeq 10^{-34}$. Evolving back the current radiation fields to the reheating era, we find $\rho_{\gamma,{\rm reh}}\simeq 10^{62}\,{\rm GeV}^{4}$ assuming instantaneous thermalization. In consistency with Turner and Widrow (1988) we find that the magnetic fields in our model after reheating, while being coherent to extremely large scales, they get diluted to such a high degree during the expansion of the universe, making it impossible for them to seed the galactic dynamo mechanism unless additional physical mechanisms are invoked Hogan (1983); Turner and Widrow (1988); Kandus et al. (2011); Subramanian (2016).